a four person committee is to be elected from an organization's membership of 11 people. how many different committees are possible
I know it is 11c4, but what is the value of this?
Thanks

Answer by jim_thompson5910(35256)

Remember that n C r = (n!)/(r!(n-r)!)


In this case, n=11 and r=4


So we get


11 C 4 = (11!)/(4!(11-4)!) = (11!)/(4!7!) = (11*10*9*8*7!)/(4!7!) = (11*10*9*8)/(4*3*2*1) = 7920/24 = 330


So 11 C 4 = 330, which means that there are 330 different committee.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim